OpenStudy (anonymous):

Integrate (e^x)/(3+e^x)^2 dx, plz i need help and id appreciate if u can explain step through step cuz i hav no idea what im doing >.<

OpenStudy (anonymous):

Try using a u-substitution where u=3+e^x :)

OpenStudy (anonymous):

you use u-substitution by letting u =3+e^x. du=e^xdx

OpenStudy (anonymous):

so u should get -(u+3)^(-1) +c

OpenStudy (anonymous):

these sorts of integral u can solve in ur head... even without using sub method

OpenStudy (anonymous):

\[u = 3+e^x \longrightarrow du=e^xdx \longrightarrow dx=e^{-x}du\] \[\int\limits \frac{e^xe^{-x}du}{u}=\int\limits \frac{du}{u^2}\] \[= -\frac{1}{u}+c\] \[=-\frac{1}{x+e^x}+c\]

OpenStudy (anonymous):

this is the full step by step solution